Emmanuel and Anthony contributed equal amounts to the purchase of a gift. Emmanuel contributed his share in five-dollar bills and Anthony gave his share in one-dollar bills. Anthony needed 12 more bills than Emmanuel. How much did each contribute toward the gift?
Answer by ankor@dixie-net.com(22740)

Emmanuel and Anthony contributed equal amounts to the purchase of a gift.
Emmanuel contributed his share in five-dollar bills and Anthony gave his share in one-dollar bills.
Anthony needed 12 more bills than Emmanuel. How much did each contribute toward the gift?
:
Let x = no. bills contributed by E
the
(x+12) = no. bills contributed by A
:
They contributed equal amts, so we have:
:
5x = 1(x+12)
5x = x + 12
5x - x = 12
4x = 12
x = 3 ea $5 bills by E
then
3 + 12 = 15 ea $1 bills by A
;
You can see they are equal $